# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 6.2. Motivation: A conjectural Large Deviation Principle
We begin by formulating a conjecture, which serves as motivation for the conjectures stated in the introduction of the paper. Recall that a sequence of measures $\Gamma_{N}$ on a topological space $\mathcal{M}$ satisfies a Large Deviation Principle (LDP) with speed $N$ and rate function $F\colon\mathcal{M}\rightarrow$$\infty$$\infty$] if $F$ is lsc and for any (Borel) measurable subset $\mathcal{B}$ of $\mathcal{M}$
@@
The following conjectures hold:
@@
$(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}\sim e^{-N\left(\beta E+\chi\right)}(\delta_{N})_{*}\left(\mu_{0}^{\otimes N}\right),$
as probability measures on $\mathcal{P}(X),$ where $\chi=0$ on $\mathcal{P}(X)_{0}$ and $\chi=\infty$ away from $\mathcal{P}(X)_{0}.$ Next, according to Sanov’s classical LDP [50]
@@
Hence, formally, $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ satisfies a LDP with rate functional $\beta F+\chi.$ Finally, since $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ is supported on $\mathcal{P}(X)_{0}$ such an LDP is equivalent to the LDP for $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ in Conjecture 6.4. The argument for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ is similar.
@@
Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 6.11. Hence, the assumed LDP for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ yields
@@
showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 6.7 is an equality. As a consequence, $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff there exists $\gamma>1$ such that $F(-\gamma)_{0}>-\infty.$ By Theorem 3.7, assuming that the Futaki character of $(X,\Delta)$ vanishes, this means that $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a KE metric, which by the solution of the YTD conjecture equivalently means that $(X,\Delta)$ is K-polystable (see [42] for the case $\Delta=0$ with $X$ smooth and [77, 79] for the case of a general log Fano variety $(X,\Delta)).$ Finally, the last statement of the proposition is shown as above.
Next, to prove that Conjecture 6.4 A implies Conjecture 1.6, assume that $(X,\Delta)$ admits a KE metric. It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 6.6 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.14 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric). As a consequence, $\int(\nu^{(N)})_{\boldsymbol{m}_{N}}$ is finite for $N$ sufficiently large. This means that $\mu_{-1}^{(N)}(=\mu_{\boldsymbol{m}_{N}}^{(N)})$ is a well-defined probability measure for $N\gg 1.$ Thus, by assumption, $(\delta_{N})_{*}\text{$\mu_{-1}^{(N)}$ }$ satisfies a LDP with rate functional $F_{-1}(\mu)-C$ on $\mathcal{P}(X)_{0}.$ By standard results for LDPs, to prove the convergence towards $\mu_{\text{KE}}$ in 1.1 it is thus enough to show that $\mu_{\text{KE}}$ is the unique minimizer of the rate functional. But this is shown in Theorem 3.7.
@@
But this implication fails, in general. Indeed, by [15, Thm 5.11], $F_{-1}$ is lsc on $\mathcal{P}(X)$ iff $X$ admits a KE metric and $\mathcal{G}$ is trivial (which, in turn, is equivalent to $X$ being K-stable [42]). Hence, already $X=\mathbb{P}^{1}$ is a counterexample. But there are also counterexamples when $\mathcal{G}$ is trivial. Indeed, $F_{-1}$ is bounded from below on $\mathcal{P}(X)$ iff $(X,\Delta)$ is K-semistable (see [75] for the case when $X$ is smooth and $\Delta$ is trivial and the proof of [1, Thm 2.5] for the general case). However, there are examples of K-semistable log Fano manifolds $(X,\Delta)$ with $\mathcal{G}$ is trivial, such that $(X,\Delta)$ is not K-stable (for example, this can be achieved when $X=\mathbb{P}^{1}$ and $\Delta$ consists of three points).
It is natural to ask about the relation between the assumptions in Conjectures A(i) and A(ii) (and likewise for B(i) and B(ii)). If inequality 6.9 can be strengthened as follows, then the second assumption implies the first:
@@
The first statement follows directly from the observation that $\int_{X^{N}}\nu_{-\gamma}^{(N)}<\infty\implies$ $\gamma<\gamma^{(N)}(X,\Delta)$ (by the openness property of local integrability thresholds). The second statement is shown in a similar way, using Prop 6.11 below. ∎
@@
As shown in [57] (see also [19] for an analytic proof) $\delta_{k}(X)\geq\gamma^{(N_{k})}(X),$ where $\delta_{k}(X)$ is the invariant introduced in [57], defined in terms of log canonical thresholds of basis divisors on $X.$ Next, in the case of a toric Fano variety which is Gorenstein (i.e. the corresponding moment polytope is reflexive) it follows from [65, Cor 2.13] that $\delta_{k}(X)<\delta(X)$ for $k$ sufficiently large, when $\delta(X)<1.$ Finally, when $n=2$ it is shown in [65, Cor 2.13] that
@@
6.3. Gibbs polystability vs finiteness of the restricted partition function $Z_{\boldsymbol{m}_{N}}$
Lemma 6.9.
Given a log Fano manifold $(X,\Delta),$ there exists a real character $\chi_{N}$ on $\mathcal{G}$ such that $g^{*}\nu_{N}=\chi_{N}(g)\nu_{N}$ for any $g\in\mathcal{G}.$ More precisely, given $W$ in the Lie algebra of $\mathcal{G}$, $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)},$ for $t\in\mathbb{C},$ where $\text{Fut}_{(X,\Delta),k}(W)$ denotes the quantized Futaki invariant at level $k.$
Proof.
By basic linear algebra, $g^{*}\nu_{N}=\left|\det A(g)\right|^{-2/k}\nu_{N}$ where $A(\cdot)$ denotes the representation of $\mathcal{G}$ on $H^{0}(X,-kK_{(X,\Delta)}).$ Hence, $\chi_{N}(g)\coloneqq\left|\det A(g)\right|^{-2/k}$ is a real character on $\mathcal{G}$ and $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)}$ by the very definition of $\text{Fut}_{(X,\Delta),k}(W).$
@@ 6.3. Gibbs polystability vs finiteness of the restricted partition function $Z_{\boldsymbol{m}_{N}}$
First, by the analytic definition of the lct (formula 2.4), $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}).$ Next, note that there exists $\epsilon_{N}\in]0,\infty[$ such that the closure of $\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ is contained in $(X^{N})_{\text{ss}}$ (since $\left\{\boldsymbol{m}_{N}=0\right\}$ is compact and contained in the open set $(X^{N})_{\text{ss}}).$ Hence, $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ implies that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right).$ To prove the converse, assume that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right)$ and consider the integral of $\nu^{(N)}$ over a neighbourhood $U$ of a given point $(x_{1},\dots,x_{N})$ in $(X^{N})_{\text{ss }}.$ By the Kempf–Ness theorem [64, Thm 3.3], a point $(x_{1},\dots,x_{N})\in(X^{N})_{\text{ss}}$ iff $\overline{\mathcal{G}(x_{1},\dots,x_{N})}$ intersects $\left\{\left\|\boldsymbol{m}_{N}\right\|=0\right\}.$ This means that there exists $g\in\mathcal{G}$ such that $g(x_{1},\dots,x_{N})\in\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ and, as a consequence, there exists a neighborhood $U$ of $(x_{1},\dots,x_{N})$ such that $gU\subset\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}.$ Thus, by assumption, $\int_{gU}\nu^{(N)}<\infty$ and since the integrability of $\nu^{(N)}$ is $\mathcal{G}$–invariant (by the previous lemma) we deduce that $\int_{U}\nu^{(N)}<\infty,$ showing that $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}),$ as desired. ∎
@@
$\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $Z_{\boldsymbol{m}_{N}}<\infty.$ In particular, $(X,\Delta)$ is Gibbs polystable iff it is Gibbs semistable and $Z_{\boldsymbol{m}_{N}}<\infty$ for all sufficiently large $N.$
@@
(6.11) $C^{-1}\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}\leq\int_{U_{\epsilon}}\nu^{(N)}\leq C\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}$
But this follows from combining the inequalities 6.12 below with the first identity in Proposition 6.14 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
6.4. Relating Gibbs polystability and the restricted partition function $Z_{\boldsymbol{m}_{N}}$ to integrals over $(X^{N})_{\mathrm{s}}/\mathcal{G}$
The main virtue of the following proposition, as compared with Prop 6.11, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
@@
Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 6.10, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small. Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 6.10. As a consequence, if $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then the integral of $\nu^{(N)}$ over $U_{\epsilon}$ is finite. It will thus be enough to show that
@@
Hence, the desired formula 6.13 follows from formula 6.14. ∎
@@
Formula 6.14 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [20], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
@@
$\int_{\left\{\boldsymbol{m}_{N}=0\right\}}f_{N}\nu_{\boldsymbol{m}_{N}}^{(N)}=\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}$
@@
We recall that 0 is a regular value of $\boldsymbol{m}_{N}$ iff the stable locus $(X^{N})_{\text{s}}$ coincides with both the semistable and the polystable locus [64, Cor 3.10]. In general, $(X^{N})_{\text{s}}$ is a Zariski open subset of the semistable locus and $(X^{N})_{\text{s}}/\mathcal{G}$ is a Zariski open subset (which as an orbifold) of the projective variety defined by the GIT quotient of $X^{N}$ under the action of $\mathcal{G}.$ Before turning to the proof of this proposition we note that, combining the previous two propositions gives the following
@@
To this end it is enough to show that the projection $p$ in 6.16 satisfies
@@
Finally, Prop 6.14 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$

### v3

@@ 6.3. Motivation: A conjectural Large Deviation Principle
We begin by formulating a conjecture, which serves as motivation for the conjectures stated in the introduction of the paper. Recall that a sequence of measures $\Gamma_{N}$ on a topological space $\mathcal{X}$ satisfies a Large Deviation Principle (LDP) with speed $N$ and rate function $F\colon\mathcal{X}\rightarrow]-\infty,\infty]$ if $F$ is lsc and for any (Borel) measurable subset $\mathcal{B}$ of $\mathcal{X}$
@@
The following conjectures hold when $N$ is a sequence of regular values:
@@
Without assuming that $N$ is a sequence of regular values the corresponding LDPs also hold on $\mathcal{P}(X)$ when $\left\{\boldsymbol{m}_{N}=0\right\}$ is replaced by $\left\{|\boldsymbol{m}_{N}|\leq\epsilon_{N}\right\}$ for any appropriate sequence $\epsilon_{N},$ with rate functional $F_{\beta}+\chi_{0},$ where $\chi_{0}=0$ on $\mathcal{P}(X)_{0}$ and $\chi_{0}=\infty$ away from $\mathcal{P}(X)_{0}.$
@@
$(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}\sim e^{-N\left(\beta E+\chi_{0}\right)}(\delta_{N})_{*}\left(\mu_{0}^{\otimes N}\right),$
as probability measures on $\mathcal{P}(X),$ where $\chi_{0}=0$ on $\mathcal{P}(X)_{0}$ and $\chi_{0}=\infty$ away from $\mathcal{P}(X)_{0}.$ Next, according to Sanov’s classical LDP [63]
@@
Hence, formally, $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ satisfies a LDP with rate functional $\beta F+\chi_{0}.$ Finally, since $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ is supported on $\mathcal{P}(X)_{0}$ such an LDP is equivalent to the LDP for $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ in Conjecture 6.7. The argument for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ is similar.
@@
Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 6.13. Hence, the assumed LDP for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ yields
@@
showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 6.9 is an equality. As a consequence, $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff there exists $\gamma>1$ such that $F(-\gamma)_{0}>-\infty.$ By Theorem 3.7, assuming that the Futaki character of $(X,\Delta)$ vanishes, this means that $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a KE metric, which by the solution of the YTD conjecture equivalently means that $(X,\Delta)$ is K-polystable (see [53] for the case $\Delta=0$ with $X$ smooth and [95, 97] for the case of a general log Fano variety $(X,\Delta)).$ Finally, the last statement of the proposition is shown as above.
Next, to prove that Conjecture 6.7 A implies Conjecture 1.8, assume that $(X,\Delta)$ admits a KE metric. It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 6.8 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.16 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric). As a consequence, $\int(\nu^{(N)})_{\boldsymbol{m}_{N}}$ is finite for $N$ sufficiently large. This means that $\mu_{-1}^{(N)}(=\mu_{0}^{(N)})$ is a well-defined probability measure for $N\gg 1.$ Thus, by assumption, $(\delta_{N})_{*}\text{$\mu_{-1}^{(N)}$ }$ satisfies a LDP with rate functional $F_{-1}(\mu)-C$ on $\mathcal{P}(X)_{0}.$ By standard results for LDPs, to prove the convergence towards $\mu_{\text{KE}}$ in 1.1 it is thus enough to show that $\mu_{\text{KE}}$ is the unique minimizer of the rate functional. But this is shown in Theorem 3.7.
@@
But this implication fails, in general. Indeed, by [16, Thm 5.11], $F_{-1}$ is lsc on $\mathcal{P}(X)$ iff $X$ admits a KE metric and $\mathcal{G}$ is trivial (which, in turn, is equivalent to $X$ being K-stable [53]). Hence, already $X=\mathbb{P}^{1}$ is a counterexample. But there are also counterexamples when $\mathcal{G}$ is trivial. Indeed, $F_{-1}$ is bounded from below on $\mathcal{P}(X)$ iff $(X,\Delta)$ is K-semistable (see [93] for the case when $X$ is smooth and $\Delta$ is trivial and the proof of [2, Thm 2.5] for the general case). However, there are examples of K-semistable log Fano manifolds $(X,\Delta)$ with $\mathcal{G}$ trivial, such that $(X,\Delta)$ is not K-stable (for example, this can be achieved when $X=\mathbb{P}^{1}$ and $\Delta$ consists of three points).
It is natural to ask about the relation between the assumptions in Conjectures A(i) and A(ii) (and likewise for B(i) and B(ii)). If inequality 6.11 can be strengthened as follows, then the second assumption implies the first:
@@
The first statement follows directly from the observation that $\int_{X^{N}}\nu_{-\gamma}^{(N)}<\infty\implies$ $\gamma<\gamma^{(N)}(X,\Delta)$ (by the openness property of local integrability thresholds). The second statement is shown in a similar way, using Prop 6.13 below. ∎
@@
As shown in [73] (see also [20] for an analytic proof) $\delta_{k}(X)\geq\gamma^{(N_{k})}(X),$ where $\delta_{k}(X)$ is the invariant introduced in [73], defined in terms of log canonical thresholds of basis divisors on $X.$ Next, in the case of a toric Fano variety which is Gorenstein (i.e. the corresponding moment polytope is reflexive) it follows from [85, Cor 2.13] that $\delta_{k}(X)<\delta(X)$ for $k$ sufficiently large, when $\delta(X)<1.$ Finally, when $n=2$ it is shown in [85, Cor 2.13] that
@@
6.4. Gibbs polystability vs finiteness of the restricted partition function $Z_{N,0}$
@@
First, by the analytic definition of the lct (formula 2.6), $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}).$ Next, note that there exists $\epsilon_{N}\in]0,\infty[$ such that the closure of $\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ is contained in $(X^{N})_{\text{ss}}$ (since $\left\{\boldsymbol{m}_{N}=0\right\}$ is compact and contained in the open set $(X^{N})_{\text{ss}}).$ Hence, $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ implies that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right).$ To prove the converse, assume that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right)$ and consider the integral of $\nu^{(N)}$ over a neighbourhood $U$ of a given point $(x_{1},\dots,x_{N})$ in $(X^{N})_{\text{ss }}.$ By the Kempf–Ness theorem [83, Thm 3.3], a point $(x_{1},\dots,x_{N})\in(X^{N})_{\text{ss}}$ iff $\overline{\mathcal{G}(x_{1},\dots,x_{N})}$ intersects $\left\{\left\|\boldsymbol{m}_{N}\right\|=0\right\}.$ This means that there exists $g\in\mathcal{G}$ such that $g(x_{1},\dots,x_{N})\in\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ and, as a consequence, there exists a neighborhood $U$ of $(x_{1},\dots,x_{N})$ such that $gU\subset\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}.$ Thus, by assumption, $\int_{gU}\nu^{(N)}<\infty$ and since the integrability of $\nu^{(N)}$ is $\mathcal{G}$–invariant (by the previous lemma) we deduce that $\int_{U}\nu^{(N)}<\infty,$ showing that $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}),$ as desired. ∎
@@
First assume that 0 is a regular value of $\boldsymbol{m}_{N}.$ Then $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $Z_{N,0}<\infty.$ In particular, $(X,\Delta)$ is Gibbs polystable iff it is Gibbs semistable and $Z_{N,0}<\infty$ for all sufficiently large $N.$ In general, $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $Z_{N,\epsilon_{N}}<\infty$ when $\epsilon_{N}$ is sufficently small. In particular, $(X,\Delta)$ is Gibbs polystable iff it is Gibbs semistable and $Z_{N,\epsilon_{N}}<\infty$ when $N$ is sufficiently large and $\epsilon_{N}$ is sufficently small.
@@
(6.13) $C^{-1}\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{0}^{(N)}\leq\int_{U_{\epsilon}}\nu^{(N)}\leq C\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{0}^{(N)}$
But this follows from combining the inequalities 6.14 below with the first identity in Proposition 6.16 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
6.5. Relating Gibbs polystability and the restricted partition function $Z_{N,0}$ to integrals over $(X^{N})_{\mathrm{s}}/\mathcal{G}$
The main virtue of the following proposition, as compared with Prop 6.13, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
@@
Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 6.12, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small. Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 6.12. As a consequence, if $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then the integral of $\nu^{(N)}$ over $U_{\epsilon}$ is finite. It will thus be enough to show that
@@
Hence, the desired formula 6.15 follows from formula 6.16. ∎
@@
Formula 6.16 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [21], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
@@
$\int_{\left\{\boldsymbol{m}_{N}=0\right\}}f_{N}\nu_{0}^{(N)}=\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}$
@@
We recall that 0 is a regular value of $\boldsymbol{m}_{N}$ iff the stable locus $(X^{N})_{\text{s}}$ coincides with both the semistable and the polystable locus [83, Cor 3.10]. In general, $(X^{N})_{\text{s}}$ is a Zariski open subset of the semistable locus and $(X^{N})_{\text{s}}/\mathcal{G}$ is a Zariski open subset (which as an orbifold) of the projective variety defined by the GIT quotient of $X^{N}$ under the action of $\mathcal{G}.$ Before turning to the proof of this proposition we note that, combining the previous two propositions gives the following
@@
To this end it is enough to show that the projection $p$ in 6.18 satisfies
@@
Finally, Prop 6.16 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$